OpenStudy (anonymous):

Find the maclaurin series of 2x/(1-3x^2)

OpenStudy (anonymous):

i was thinking i could just change it to 2x * 1/(1-3x^2) but its not working

OpenStudy (anonymous):

multiply both sides by 2x to get ur answer

OpenStudy (anonymous):

does the xsqrt3 come from the 3x^2?

OpenStudy (anonymous):

wait i made a mistake

OpenStudy (anonymous):

if \( |3x^2|<1 \) then u can expand \( \frac{1}{1-3x^2} \) like this \( \frac{1}{1-3x^2}=1+3x^2+9x^4+... \) am i right?

OpenStudy (anonymous):

think of \( \frac{1}{1-x}=1+x+x^2+... \) for \( |x|<1 \) and replace x with 3x^2

OpenStudy (anonymous):

then u just need to multiply it by 2x to get the right answer

OpenStudy (anonymous):

I get that and yeah I get SUM(3x^2)^n but if it is a maclaurin series doesn't x = 0 and i just get all 0's ?

OpenStudy (anonymous):

or is it just that I am just looking for the series and it doesn't matter. just to get \[2x \sum_{n=0}^{\inf} (3x^2)^n\]

OpenStudy (anonymous):

it doesn't matter i think

OpenStudy (anonymous):

Your function is indeed equal to zero at the orgin.